OpenStudy (anonymous):

If ‘ abca ’ is a 4-digit number which equals to a^b x c^a . Find a, b and c

OpenStudy (anonymous):

1000a+100b+10c+a = a^b * c^a 1001a+100b+10c = a^b * c^a

OpenStudy (anonymous):

i think it should be abcd not abca

OpenStudy (anonymous):

wait i got it

OpenStudy (anonymous):

LHS is divisible by 'a', so RHS must also be divisible by 'a'. Therefore, let (100b + 10c) / a = J (where J is an integer), so, [2] 10(10b + c) = aJ The factors of this LHS include 1, 2, 5, 10, (10b + c) and any factors of (10b + c). If 'a' is a single digit, it can only be 2, 5, or a factor of (10b + c). (note that 'a' cannot be 1) I'm going to dismiss the factors of (10b + c) as I haven't a clue what they are. First, let a = 5. Therefore, from [2] we get : 100b + 10c = 5J Plugging these into [1] gives : (5^b)(c^5) = 5005 + 5J Dividing through by 5 : [3] [5^(b-1)](c^5) = 1001 + J Now, 100b + 10c = 5J is a minimum of 110 if b = 1, c = 1, and is a maximum of 990 if b = 9, c = 9. Thus, 110 <= 5J <= 990, or, 22 <= J <= 198. So in [3], 1023 <= [5^(b-1)](c^5) <= 1199 Testing shows that c cannot be 1, 2, 3 or >4, but it can be true for c = 4, giving b = 1, but, [1] is then not true. Therefore, 'a' is not equal to 5. So try a = 2. In [1], we have : (2^b)(c^2) = 2002 + 100b + 10c Dividing by 2 : [2^(b-1)](c^2) = 1001 + 50b + 5c Minimum value for 1001 + 50b + 5c is 1056. Maximum value is 1496. So, 1056 <= [2^(b-1)](c^2) <= 1496 Because of c^2, we must be able to take the square root. Therefore, 32.5 <= {2^[(b-1)/2]}c <= 38.7 Trying values for c, it's found that only 2^2 * 9 = 36, works. Thus, c = 9 and (b-1)/2 = 2, so b = 5.

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):

BUT 2592 is not equal to 2^5 * 9^2

OpenStudy (anonymous):

i m not sure

OpenStudy (anonymous):

2592 is equal to (2^5)*(9^2).\[2^5\cdot 9^2=32\cdot 81 = 2592\]

OpenStudy (anonymous):

oh yes.... thought 9^2 =18

OpenStudy (anonymous):

my answer is correct..i was in doubt it is wrong solved by me. thank u so much

OpenStudy (anonymous):

Thank u so much @sauravshakya and @joemath314159